One angle of a triangle measures 72 degrees more than the smallest, while a third angle is twice the smallest. Find the measure of each angle.
Answer by addingup(3677)

The angles of EVERY triangle add up to 180 degrees.
Let the angles be A, B, and C
A = B+72
C = 2B
~~~~~~~~~~~~~~~~~~~~~
B+72+B+2B = 180
4B = 108
B = 27
A = 27+72 = 99
C = 2*27 = 54
Happy learning!